OpenStudy (anonymous):

HELPP!!! Some one please let me understand how to do this and the formula Find the 25th term of the following sequence22, 26, 30... thank you

OpenStudy (goldeverything):

118

OpenStudy (anonymous):

help me understand how you got there please @GoldEverything

OpenStudy (goldeverything):

well your looking for the 25th sequence correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

isn't there like a formula?

OpenStudy (goldeverything):

not really if you look at the number you had 22,26, and 30

OpenStudy (goldeverything):

you can see a pattern of plus 4

OpenStudy (goldeverything):

so i started at 22 and added 4 25 times untill i got 118

OpenStudy (anonymous):

ooh ok then thank you i get it now

OpenStudy (goldeverything):

anytime

OpenStudy (mathstudent55):

You can find the answer by adding 4 until you get tot he 25th term, or you can use this formula. \(a_n = a_1 + (n - 1)d\) where \(a_1\) = first term n = number of the term you need d = common difference In your case, do: \(a_{25} = 22 + ( 25 - 1)4\) \(a_{25} = 22 + (24)4\) \(a_{25} = 22 + 96\) \(a_{25} = 118\)

OpenStudy (goldeverything):

so basically i did it right just a easier way? @mathstudent55

OpenStudy (mathstudent55):

Yes, you did it an easier way. In this case your way was relatively fast. The problem is if you were asked for the 1000th term. However, you were not correct when asked if there is a formula, and you answered "not really." In fact, as you can see, there is a formula. Very often in math, there are several ways of solving a problem. Some could be more work, and some could be less work. The important thing is that they all be valid methods that will give you the correct solution. In this case, both of our methods are valid, and they lead to the same correct solution.